OpenStudy (anonymous):

so steps using zero property solve this equation pls: y^3=-22y^2-40y

OpenStudy (jhannybean):

First step is to move everything to the left side of your equation. What would you get?

OpenStudy (jhannybean):

You can do this by adding +22y^2 +40y to both sides.

OpenStudy (anonymous):

if i subtract y^3 i get 0=-23y-40y right

OpenStudy (jhannybean):

You would subtract y^3 from both sides then, but it's easier to make everything positive than keep it negative.

OpenStudy (anonymous):

so what i get after subtract y^3

OpenStudy (jhannybean):

-y^3 + y^3 = -y^3 -22y^2 -40y What would this simplify to?

OpenStudy (anonymous):

but y^3 is not a square root how do i solve for y?

OpenStudy (jhannybean):

Follow me on this. It'll make sense one we work it step by step. Did you simplify the expression i presented in my last post?

OpenStudy (anonymous):

simply ?

OpenStudy (anonymous):

so i devided by y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

dividing by y will not help us.